I want to make a mixture of feed that is 16% protein. I need a total of 2000 pounds, so I'm after 320# of protein.
I can mix soybean which is 48.5% and milo which is 9%. I'm not sure how to figure out the answer.

Answer by richwmiller(17219)

We need a total of 2000 lb. at 16%
We want to know how many lb. at 48.5% to mix with 9%
a+b=2000,
0.485*a+0.09*b=0.16*2000
a=2000-b
0.485*(2000-b)+0.09*b=320
970.0-0.485b+0.09*b=320
-0.395*b=-650.0
b=1645.56962
a=2000-b
a=354.43038 lb. at 48.5%
b=1645.56962 lb. at 9%
check
0.485*354.43038+0.09*1645.56962=0.16*2000
171.898734+148.101266=320
320.0=320
ok
codemx1